OpenStudy (lgbasallote):

\(\large \mathbf{\color{maroon}{L} \color{violet}{G} \color{orange}{B} \color{darkblue}{A} \color{gold}{R} \color{brown}{I} \color{pink}{D} \color{purple}{D} \color{green}{L} \color{white}{E}}\) \(\LARGE \int \frac{1}{\sqrt x - \sqrt[3]{x}}dx\) Hint: Be Creative!

OpenStudy (lgbasallote):

@Ishaan94

OpenStudy (lgbasallote):

this is very easy once you found the first step :P

OpenStudy (mimi_x3):

hint please; i don't want to go on the track...what substitution?

OpenStudy (lgbasallote):

if i say the substitution then it's already solved :P that's the key haha

hero (hero):

not hard

OpenStudy (lgbasallote):

shhh hero :S

hero (hero):

I didn't say anything relevant :P

OpenStudy (lgbasallote):

i assume you found the substitution already..

hero (hero):

I'm not saying anything :P

OpenStudy (mimi_x3):

let u =x^(1/3) ?

OpenStudy (lgbasallote):

haha no mimi ^_^

OpenStudy (mimi_x3):

well, times the denominator and numerator by something? i dont want to go on the wrong track it will get messy

OpenStudy (lgbasallote):

nope either :) just a simple u substitution

OpenStudy (mimi_x3):

u= x^(1/2) ?

OpenStudy (lgbasallote):

still no :p haha

OpenStudy (mimi_x3):

u =x^(1/6) ?

OpenStudy (lgbasallote):

yup! yay :D now solve it >:))

OpenStudy (mimi_x3):

lol, then its long division too lazy

OpenStudy (lgbasallote):

hahaha =)) that's why it's an lgbariddle ;D

OpenStudy (mimi_x3):

Let, Hero do it.. :P

OpenStudy (lgbasallote):

hmmm..

OpenStudy (anonymous):

Take u =x^(1/6) and then go ahead with partial fraction . Tiring :P

OpenStudy (blockcolder):

Yeah. That's the only way to do it (I think).

OpenStudy (lgbasallote):

ahh the beauty of lgbariddles >:))

OpenStudy (anonymous):

No it is plum \[\int\limits_{}^{} \large \frac{6u^5} {u^2(u-1)}\] \[\large 6\int\limits_{}^{}\frac{u^3}{u-1} = \large 6\int\limits_{}^{}\frac{u^3-1^3}{u-1} +6\int\limits_{}^{}\frac{1}{u-1}\] Now continue :D

OpenStudy (anonymous):

Answer should come within 2 steps :D

OpenStudy (wasiqss):

shivam you have nnothing now :P

OpenStudy (anonymous):

lol. I forgot du everywhere :P

OpenStudy (wasiqss):

lol du= whatever :P

OpenStudy (anonymous):

"shivam you have nnothing now :P "-->????

OpenStudy (blockcolder):

Is the first integral an arctan of something?

OpenStudy (anonymous):

You got to be kidding me @blockcolder just apply u^3-1^3 = (u-1)(u^2+1+u) and cancel the numerator u-1 and denominator u-1

OpenStudy (blockcolder):

And then complete the square of the denominator and viola! A wild arctan appears!

OpenStudy (anonymous):

@blockcolder \[\large 6\int\limits\limits_{}^{}{(u^2+1+u)}du +6\int\limits\limits_{}^{}\frac{1}{u-1}\] \[\large 6u^3/3 + 6u+ 3u^2 + 6\log(u-1)\]

OpenStudy (anonymous):

substitute u= x^(1/6) and voila :P

OpenStudy (blockcolder):

Oh, right. I thought the u^3-1 is in the denominator. Guess I should clean my glasses.

OpenStudy (anonymous):

LOL

OpenStudy (blockcolder):

Yeah, that happens a lot with me. =))